# Phase-field modeling of radiation-induced composition redistribution: An application to additively manufactured austenitic Fe-Cr-Ni

### v1

@@ A.1 Sink strength of GB
$J_{\nu}.\hat{n}=D_{\nu}\alpha^{-1}(c_{\nu}-c^{e}_{\nu}),$ (A1)
where $J_{\nu}.\hat{n}$ is the flux of point defect normal to the GB plane. In the PF implementation, the total absorption rate across a planar GB per unit area can be written as:
@@
The absorption rates between the sharp-interface model with RBC and the PF model can be related as $2J_{\nu}.\hat{n}=S_{\nu}$. However, since $c_{\nu}$ is expected to vary spatially within the GB, one cannot accurately evaluate the above integral. As an approximation, we evaluated it by assuming that $c_{\nu}$ is uniform within the diffuse GB, and related the coefficients between the sharp-interface and PF models as:
@@ A.2 Steady-state concentrations
$\frac{\partial c_{\nu}}{\partial t}=-\frac{dJ_{x,\nu}}{dx}+P_{\nu}-k^{2}_{b,\nu}D_{\nu}(c_{\nu}-c^{e}_{\nu}),$ (A5)
with an RBC (Eq. A1) applied to model the GB sink. Here, recombination between vacancies and SIAs, which is only expected to contribute at low temperatures, has been neglected. At steady state ($\partial c_{\nu}/\partial t=0$), the point defect concentrations at the grain center ($dJ_{x,\nu}/dx=0$) are given by:
$c_{\nu}=c^{e}_{\nu}+\frac{P_{\nu}}{k^{2}_{b,\nu}D_{\nu}}.$ (A6)
Furthermore, neglecting bulk sinks ($k^{2}_{b,\nu}=0$) and integrating Eq. A5 from the grain center to the GB plane gives us:
@@
where $d$ is the grain length and $J^{g}_{x,\nu}(x=0)$ and $J^{b}_{x,\nu}(x=\pm\frac{d}{2})$ are the fluxes at the GB plane and grain center, respectively. In the RBC model, $J_{x}$ varies monotonically from the grain center to the GB plane, with the maximum in $J_{x}$ occurring at the GB plane. Noting that $J^{b}_{x}=0$, and $J^{g}_{x,\nu}(x=0)$ is given by Eq. A1, we obtain the concentration at the GB, as:
$c_{\nu}=c^{e}_{\nu}+\frac{P\,d}{2D_{\nu}\alpha^{-1}}.$ (A8)
@@
$c_{\nu}=c^{e}_{\nu}+\frac{P\,d}{\delta\,C_{\text{sink}}k^{2}_{g}D_{\nu}}.$ (A9)
@@ ACKNOWLEDGMENTS
Using Onsager transport and reciprocity relations, the partial fluxes of atomic elements $J^{\nu}_{\phi}$ are given in terms of the chemical potential gradients of atomic elements $\mu_{k}$ and point defects $\mu_{\nu}$ as:
$\boldsymbol{J}^{\nu}_{\phi}$ $=-\sum^{K}_{k=1}L^{\nu}_{\phi k}\nabla\left[\mu_{k}+\text{sign}(\nu)\mu_{\nu}\right],$ (S1.1a)
where $\text{sign}(V)=-1$ and $\text{sign}(I)=+1$. The point defect fluxes are given by $\boldsymbol{J}_{\nu}=\text{sign}(\nu)\sum^{K}_{k=1}\boldsymbol{J}^{\nu}_{k}$. By using the Gibbs-Duhem relation $\sum^{K}_{k=1}c_{k}\nabla\mu_{k}=0$, and substituting $c_{1}\approx 1-\sum^{K}_{k=2}c_{k}$, we can write $\nabla\mu_{1}=-\sum^{K}_{k=2}c_{k}\nabla(\mu_{k}-\mu_{1})$ and $\nabla\mu_{k(\neq 1)}=(1-\sum^{K}_{k=2}c_{k})\nabla(\mu_{k}-\mu_{1})+\sum^{K}_{j=2}c_{j}\nabla(\mu_{k}-\mu_{j})$. The independent fluxes ($\phi=2:K$) can then be written as:
$\boldsymbol{J}^{\nu}_{\phi}=-\sum^{K}_{k=2}\left(L^{\nu}_{\phi k}-c_{k}\sum^{K}_{j=1}L^{\nu}_{\phi j}\right)\nabla(\mu_{k}-\mu_{1})$
$\hskip 42.67912pt-\sum_{\nu=V,I}\left(\sum^{K}_{j=1}\text{sign}(\nu)L^{\nu}_{\phi j}\right)\nabla\mu_{\nu}$ (S1.2a)
$\boldsymbol{J}_{\nu}=-\sum^{K}_{k=2}\left[\text{sign}(\nu)\left(\sum^{K}_{j=1}L^{\nu}_{kj}-c_{k}\sum^{K}_{i=1}\sum^{K}_{j=1}L^{\nu}_{ij}\right)\right]\nabla(\mu_{k}-\mu_{1})$
$\hskip 42.67912pt-\left(\sum^{K}_{i=1}\sum^{K}_{j=1}L^{\nu}_{ij}\right)\nabla\mu_{\nu}.$ (S1.2b)
The total atomic fluxes are $\boldsymbol{J}_{\phi}=\sum_{\nu=V,I}\boldsymbol{J}^{\nu}_{\phi}$. By defining $L_{\phi\nu}=\sum^{K}_{j=1}\text{sign}(\nu)L^{\nu}_{\phi j}$, $L_{\nu\nu}=\sum^{K}_{i=1}\sum^{K}_{j=1}L^{\nu}_{ij}$, $L^{1}_{\nu k}=\text{sign}(\nu)\left(\sum^{K}_{j=1}L^{\nu}_{kj}-c_{k}L_{\nu\nu}\right)$, and $L^{1}_{\phi k}=\sum_{\nu}L^{\nu}_{\phi k}-\,c_{k}\sum_{\nu}\text{sign}(\nu)L_{\phi\nu}$, we obtain the total atomic and point defect fluxes as:
$\boldsymbol{J}_{\phi}$ $=-\sum^{K}_{k=2}L^{1}_{\phi k}\nabla\mu_{k1}-\sum_{\nu=V,I}L_{\phi\nu}\nabla\mu_{\nu},$ (S1.3a)
$\boldsymbol{J}_{\nu}$ $=-\sum^{K}_{k=2}L^{1}_{\nu k}\nabla\mu_{k1}-L_{\nu\nu}\nabla\mu_{\nu}.$ (S1.3b)
@@
Appendix S4 Density-based CALPHAD free energy
The free energy density ($f_{\text{C}}$) is described as:
$G^{ideal}_{mix}=RT(c_{Fe}\ln{c_{Fe}}+c_{Cr}\ln{c_{Cr}}+c_{Ni}\ln{c_{Ni}}).$ (S4.2)
${}^{o}G_{\phi}={}^{o}H^{fcc}_{\phi}\sigma^{2}-T\,{}^{o}S^{fcc}_{\phi}\sigma.$ (S4.3)
And the molar excess Gibbs energy of mixing is given by:
$+c_{Fe}c_{Cr}c_{Ni}L^{fcc}_{Fe,Cr,Ni}\sigma^{2}.$ (S4.4)
${}^{o}H^{fcc}_{Fe}=24.664\,T+154717\,T^{-1}+3.75752\times 10^{-3}\,T^{2}$
$\hskip 42.67912pt+1.178538\times 10^{-7}\,T^{3}-236.7,$ (S4.5a)
${}^{o}H^{fcc}_{Cr}=26.908\,T+278500\,T^{-1}-1.89435\times 10^{-3}\,T^{2}$
$\hskip 42.67912pt+2.95442\times 10^{-6}\,T^{3}-1572.94,$ (S4.5b)
${}^{o}H^{fcc}_{Ni}=4.8407\times 10^{-3}\,T^{2}+22.096\,T-5179.159$ (S4.5c)
and
${}^{o}S^{fcc}_{Fe}=7.51504\times 10^{-3}\,T+24.664\,\ln{T}+77358.5\,T^{-2}$
$\hskip 42.67912pt+1.767807\times 10^{-7}\,T^{2}-107.7517,$ (S4.6a)
${}^{o}S^{fcc}_{Cr}=26.908\,\ln{T}-3.7887\times 10^{-3}\,T+139250\,T^{-2}$
$\hskip 42.67912pt+4.43163\times 10^{-6}\,T^{2}-130.735,$ (S4.6b)
${}^{o}S^{fcc}_{Ni}=9.6814\times 10^{-3}\,T+22.096\,\ln{T}-95.758.$ (S4.6c)
$L^{fcc}_{Fe,Cr}=(10833-7.477)+(-1410)(c_{Fe}-c_{Cr}),$ (S4.7a)
$L^{fcc}_{Fe,Ni}=(-12054+3.274\,T)+(11082-4.45\,T)(c_{Fe}-c_{Ni})$
$\hskip 42.67912pt+(-726)(c_{Fe}-c_{Ni})^{2},$ (S4.7b)
$L^{fcc}_{Cr,Ni}=(8030-12.88\,T)+(33080-16.036\,T)(c_{Cr}-c_{Ni})$ (S4.7c)
$L^{fcc}_{Fe,Cr,Ni}=(-6500)c_{Fe}+(10000-10\,T)c_{Cr}+(48000)c_{Ni}.$ (S4.7d)
@@
where $\Delta\omega^{e}_{g}=\omega^{e}_{g}-\omega^{e}_{b}$. Here, “$e$” denotes that the quantities are evaluated for the condition of a stationary GB, $\partial\eta_{i}/\partial t=0$, and a constant diffusion potential $\mu_{k1}$ across the system. For a large system, the equilibrium concentration in the bulk can be assumed to be close to nominal. Thus, considering “$e$” and $\circ$” (determined in Sec. 3.4) to be equivalent gives us $\Delta\omega^{e}_{g}=f^{g,o}-f^{b,o}-\mu^{o}_{CrFe}(c^{g,\circ}_{Cr}-c^{b,\circ}_{Cr})-\mu^{o}_{NiFe}(c^{g,\circ}_{Ni}-c^{b,\circ}_{Ni})$. At 500 C and $\sigma=0.8$, this yields $\Delta\omega^{e}_{g}=2.5\times 10^{8}$ J/m3. Thus, for a GB energy of $\gamma=1$ J/m2, $m_{0}=0$ could be set to uniquely determine $\kappa=2.2\times 10^{-9}$ J/m and $\delta=2.9$ nm by using the above relations. However, the temperature dependence of $\gamma$ of $\Delta\omega^{e}_{g}$ would lead to significant GB width variation with temperature. Moreover, $\Delta\omega^{e}_{g}$ tends to vanish as $\sigma$ approaches unity (i.e., as the free energy of the GB approaches that of the bulk), yielding unphysical widths. Therefore, to model GBs without TS, a finite $m_{0}$ must be determined by setting a value for $\delta$ per the convention in PF models. To simplify the parameterization and interpretation of results, we omit $\omega_{g}-\omega_{b}$ from Eq. 13 and set $\Delta\omega^{e}_{g}=0$ in Eqs. S5.1 and S5.2. We assume $\gamma=1$ J/m2 and $\delta=1$nm, thus yielding the parameters $m_{0}=7.5\times 10^{8}$ J/m3 and $\kappa=7.5\times 10^{-10}$ J/m.
Appendix S6 Solute-SIA binding

### v2

@@ A.1 Sink strength of GB
$J_{{\upsilon}}{\cdot}\hat{n}=D_{{\upsilon}}\alpha^{-1}(c_{{\upsilon}}-c^{e}_{{\upsilon}}),$ (A1)
where $J_{{\upsilon}}{\cdot}\hat{n}$ is the flux of point defect normal to the GB plane and $\alpha^{-1}$ is a parameter governing the rate of point defect absorption or emission by a non-ideal GB [65]. In the PF implementation, the total absorption rate across a planar GB per unit area can be written as:
@@
The absorption rates between the sharp-interface model with RBC and the PF model can be related as $2J_{{\upsilon}}{\cdot}\hat{n}=S_{{\upsilon}}$. However, since $c_{{\upsilon}}$ is expected to vary spatially within the GB, one cannot accurately evaluate the above integral. As an approximation, we evaluated it by assuming that $c_{{\upsilon}}$ is uniform within the diffuse GB, and related the coefficients between the sharp-interface and PF models as:
@@ A.2 Steady-state concentrations
$\frac{\partial c_{{\upsilon}}}{\partial t}=-\frac{dJ_{x,{{\upsilon}}}}{dx}+P_{{\upsilon}}-k^{2}_{b,{{\upsilon}}}D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}}),$ (A5)
with an RBC (Eq. A1) applied to model the GB sink. Here, recombination between vacancies and SIAs, which is only expected to contribute at low temperatures, has been neglected. At steady state ($\partial c_{{\upsilon}}/\partial t=0$), the point defect concentrations at the grain center ($dJ_{x,{{\upsilon}}}/dx=0$) are given by:
$c_{{\upsilon}}=c^{e}_{{\upsilon}}+\frac{P_{{\upsilon}}}{k^{2}_{b,{{\upsilon}}}D_{{\upsilon}}}.$ (A6)
Furthermore, neglecting bulk sinks ($k^{2}_{b,{{\upsilon}}}=0$) and integrating Eq. A5 from the grain center to the GB plane gives us:
@@
where $d$ is the grain length and $J^{g}_{x,{{\upsilon}}}(x=0)$ and $J^{b}_{x,{{\upsilon}}}(x=\pm\frac{d}{2})$ are the fluxes at the GB plane and grain center, respectively. In the RBC model, $J_{x}$ varies monotonically from the grain center to the GB plane, with the maximum in $J_{x}$ occurring at the GB plane. Noting that $J^{b}_{x}=0$, and $J^{g}_{x,{{\upsilon}}}(x=0)$ is given by Eq. A1, we obtain the concentration at the GB, as:
$c_{{\upsilon}}=c^{e}_{{\upsilon}}+\frac{P\,d}{2D_{{\upsilon}}\alpha^{-1}}.$ (A8)
@@
$c_{{\upsilon}}=c^{e}_{{\upsilon}}+\frac{P\,d}{\delta\,C_{\text{sink}}k^{2}_{g}D_{{\upsilon}}}.$ (A9)
@@ ACKNOWLEDGMENTS
Using Onsager transport and reciprocity relations, the partial fluxes of atomic elements $J^{{\upsilon}}_{\phi}$ are given in terms of the chemical potential gradients of atomic elements $\mu_{k}$ and point defects $\mu_{{\upsilon}}$ as:
$\boldsymbol{J}^{{\upsilon}}_{\phi}$ $=-\sum^{K}_{k=1}L^{{\upsilon}}_{\phi k}\nabla\left[\mu_{k}+\text{sign}({{\upsilon}})\mu_{{\upsilon}}\right],$ (S1.1a)
where $\text{sign}(V)=-1$ and $\text{sign}(I)=+1$. The point defect fluxes are given by $\boldsymbol{J}_{{\upsilon}}=\text{sign}({{\upsilon}})\sum^{K}_{k=1}\boldsymbol{J}^{{\upsilon}}_{k}$. By using the Gibbs-Duhem relation $\sum^{K}_{k=1}c_{k}\nabla\mu_{k}=0$, and substituting $c_{1}\approx 1-\sum^{K}_{k=2}c_{k}$, we can write $\nabla\mu_{1}=-\sum^{K}_{k=2}c_{k}\nabla(\mu_{k}-\mu_{1})$ and $\nabla\mu_{k(\neq 1)}=(1-\sum^{K}_{k=2}c_{k})\nabla(\mu_{k}-\mu_{1})+\sum^{K}_{j=2}c_{j}\nabla(\mu_{k}-\mu_{j})$. The independent fluxes ($\phi=2:K$) can then be written as:
$\boldsymbol{J}^{{\upsilon}}_{\phi}=-\sum^{K}_{k=2}\left(L^{{\upsilon}}_{\phi k}-c_{k}\sum^{K}_{j=1}L^{{\upsilon}}_{\phi j}\right)\nabla(\mu_{k}-\mu_{1})$
$\hskip 42.67912pt-\sum_{{{\upsilon}}=V,I}\left(\sum^{K}_{j=1}\text{sign}({{\upsilon}})L^{{\upsilon}}_{\phi j}\right)\nabla\mu_{{\upsilon}}$ (S1.2a)
$\boldsymbol{J}_{{\upsilon}}=-\sum^{K}_{k=2}\left[\text{sign}({{\upsilon}})\left(\sum^{K}_{j=1}L^{{\upsilon}}_{kj}-c_{k}\sum^{K}_{i=1}\sum^{K}_{j=1}L^{{\upsilon}}_{ij}\right)\right]\nabla(\mu_{k}-\mu_{1})$
$\hskip 42.67912pt-\left(\sum^{K}_{i=1}\sum^{K}_{j=1}L^{{\upsilon}}_{ij}\right)\nabla\mu_{{\upsilon}}.$ (S1.2b)
The total atomic fluxes are $\boldsymbol{J}_{\phi}=\sum_{{{\upsilon}}=V,I}\boldsymbol{J}^{{\upsilon}}_{\phi}$. By defining $L_{\phi{{\upsilon}}}=\sum^{K}_{j=1}\text{sign}({{\upsilon}})L^{{\upsilon}}_{\phi j}$, $L_{{{\upsilon}}{{\upsilon}}}=\sum^{K}_{i=1}\sum^{K}_{j=1}L^{{\upsilon}}_{ij}$, $L^{1}_{{{\upsilon}}k}=\text{sign}({{\upsilon}})\left(\sum^{K}_{j=1}L^{{\upsilon}}_{kj}-c_{k}L_{{{\upsilon}}{{\upsilon}}}\right)$, and $L^{1}_{\phi k}=\sum_{{{\upsilon}}}L^{{\upsilon}}_{\phi k}-\,c_{k}\sum_{{{\upsilon}}}\text{sign}({{\upsilon}})L_{\phi{{\upsilon}}}$, we obtain the total atomic and point defect fluxes as:
$\boldsymbol{J}_{\phi}$ $=-\sum^{K}_{k=2}L^{1}_{\phi k}\nabla\mu_{k1}-\sum_{{{\upsilon}}=V,I}L_{\phi{{\upsilon}}}\nabla\mu_{{\upsilon}},$ (S1.3a)
$\boldsymbol{J}_{{{\upsilon}}}$ $=-\sum^{K}_{k=2}L^{1}_{{{\upsilon}}k}\nabla\mu_{k1}-L_{{{\upsilon}}{{\upsilon}}}\nabla\mu_{{\upsilon}}.$ (S1.3b)
@@
Appendix S2 Density-based CALPHAD free energy
The free energy density ($f_{\text{C}}$) is described as:
$G^{ideal}_{mix}=RT(c_{Fe}\ln{c_{Fe}}+c_{Cr}\ln{c_{Cr}}+c_{Ni}\ln{c_{Ni}}).$ (S2.2)
${}^{o}G_{\phi}={}^{o}H^{fcc}_{\phi}\sigma^{2}-T\,{}^{o}S^{fcc}_{\phi}\sigma.$ (S2.3)
And the molar excess Gibbs energy of mixing is given by:
$+c_{Fe}c_{Cr}c_{Ni}L^{fcc}_{Fe,Cr,Ni}\sigma^{2}.$ (S2.4)
${}^{o}H^{fcc}_{Fe}=24.664\,T+154717\,T^{-1}+3.75752\times 10^{-3}\,T^{2}$
$\hskip 42.67912pt+1.178538\times 10^{-7}\,T^{3}-236.7,$ (S2.5a)
${}^{o}H^{fcc}_{Cr}=26.908\,T+278500\,T^{-1}-1.89435\times 10^{-3}\,T^{2}$
$\hskip 42.67912pt+2.95442\times 10^{-6}\,T^{3}-1572.94,$ (S2.5b)
${}^{o}H^{fcc}_{Ni}=4.8407\times 10^{-3}\,T^{2}+22.096\,T-5179.159$ (S2.5c)
and
${}^{o}S^{fcc}_{Fe}=7.51504\times 10^{-3}\,T+24.664\,\ln{T}+77358.5\,T^{-2}$
$\hskip 42.67912pt+1.767807\times 10^{-7}\,T^{2}-107.7517,$ (S2.6a)
${}^{o}S^{fcc}_{Cr}=26.908\,\ln{T}-3.7887\times 10^{-3}\,T+139250\,T^{-2}$
$\hskip 42.67912pt+4.43163\times 10^{-6}\,T^{2}-130.735,$ (S2.6b)
${}^{o}S^{fcc}_{Ni}=9.6814\times 10^{-3}\,T+22.096\,\ln{T}-95.758.$ (S2.6c)
$L^{fcc}_{Fe,Cr}=(10833-7.477)+(-1410)(c_{Fe}-c_{Cr}),$ (S2.7a)
$L^{fcc}_{Fe,Ni}=(-12054+3.274\,T)+(11082-4.45\,T)(c_{Fe}-c_{Ni})$
$\hskip 42.67912pt+(-726)(c_{Fe}-c_{Ni})^{2},$ (S2.7b)
$L^{fcc}_{Cr,Ni}=(8030-12.88\,T)+(33080-16.036\,T)(c_{Cr}-c_{Ni})$ (S2.7c)
$L^{fcc}_{Fe,Cr,Ni}=(-6500)c_{Fe}+(10000-10\,T)c_{Cr}+(48000)c_{Ni}.$ (S2.7d)
@@
where $\Delta\omega^{e}_{g}=\omega^{e}_{g}-\omega^{e}_{b}$. Here, “$e$” denotes that the quantities are evaluated for the condition of a stationary GB, $\partial\eta_{i}/\partial t=0$, and a constant diffusion potential $\mu_{k1}$ across the system. For a large system, the equilibrium concentration in the bulk can be assumed to be close to nominal. Thus, considering “$e$” and $\circ$” (determined in Sec. 3.4) to be equivalent gives us $\Delta\omega^{e}_{g}=f^{g,o}-f^{b,o}-\mu^{o}_{CrFe}(c^{g,\circ}_{Cr}-c^{b,\circ}_{Cr})-\mu^{o}_{NiFe}(c^{g,\circ}_{Ni}-c^{b,\circ}_{Ni})$. At 500 C and $\sigma=0.8$, this yields $\Delta\omega^{e}_{g}=2.5\times 10^{8}$ J/m3. Thus, for a GB energy of $\gamma=1$ J/m2, $m_{0}=0$ could be set to uniquely determine $\kappa=2.2\times 10^{-9}$ J/m and $\delta=2.9$ nm by using the above relations. However, the temperature dependence of $\gamma$ of $\Delta\omega^{e}_{g}$ would lead to significant GB width variation with temperature. Moreover, $\Delta\omega^{e}_{g}$ tends to vanish as $\sigma$ approaches unity (i.e., as the free energy of the GB approaches that of the bulk), yielding unphysical widths. Therefore, to model GBs without TS, a finite $m_{0}$ must be determined by setting a value for $\delta$ per the convention in PF models. To simplify the parameterization and interpretation of results, we omit $\omega_{g}-\omega_{b}$ from Eq. 13 and set $\Delta\omega^{e}_{g}=0$ in Eqs. S6.1 and S6.2. We assume $\gamma=1$ J/m2 and $\delta=1$nm, thus yielding the parameters $m_{0}=7.5\times 10^{8}$ J/m3 and $\kappa=7.5\times 10^{-10}$ J/m.
Appendix S7 Results of testing the adaptive time stepper
Here, we present the results of convergence study of 1D RIS simulations performed for different parameters of the time stepping scheme, IterationAdaptiveDT, in the MOOSE framework. The time stepper increases or decreases the time step to maintain a certain number of nonlinear iterations, as specified by the parameter optimal_iterations and iteration_window. For the simulations in this work, we set optimal_iterations = 8 and iteration_window = 2 as the default. As high values of optimal_iterations can result in greater time steps that could potentially increase discretization error of the simulation results, we performed a simulation for a lower value of optimal_iterations = 4 in Fig. S2. A simulation was also performed with optimal_iterations = 8 by cutting the time step sizes (dt) realized by the default simulation by a factor of 10. Another simulation was performed by limiting the maximum time step size (dt_max) to $10^{-3}$ days. The results for atomic and point defect concentrations do not change significantly for the different parameters values of the time stepper tested here; the error between the concentrations at any time is less than 1%.
Figure S2: Results from different settings in MOOSE’s IterationAdaptiveDT time stepper.(a) Simulation time step versus time. Evolution of (b) Ni and Cr RIS at GB. Evolution of vacancy (vac.) and SIA concentrations at (c) bulk and (d) GB.
Appendix S8 Supplementary results on atomic and point defect concentrations
Figure S3: Steady-state point defect concentrations near the GB in a 1D system of 1 $\mu$m grain size when under irradiation at $2\times 10^{-6}$ dpa/s and 500°C. Profiles corresponding to the RIS (Fig. 3(a)), RETS (Fig. 3(b)) and RIS+RETS (Fig. 4(a)) mechanisms are nearly identical.
Figure S4: Low magnification plot of segregation profiles from Fig. 4 corresponding to RIS and TS mechanisms in a 1D system of 1 $\mu$m grain size when under irradiation at $2\times 10^{-6}$ dpa/s and 500°C. (a) $\sigma=0.8$ and (b) $\sigma=0.9$. The non-monotonic segregation at the GB is not discernible for Ni in (a) and for Cr and Ni in (b).
Figure S5: Segregation due to very weak TS modeled by $\sigma=0.95$ in a 1D system of 1 $\mu$m grain size when under irradiation at $2\times 10^{-6}$ dpa/s and 500°C. (a) The RETS mechanism. (b) The combined RIS and TS mechanisms.
Figure S6: Point defect concentrations as a function of grain size from 1D PF simulations of RIS in Fig. 9 with bulk dislocation density $\rho_{b}=10^{14}$ m-2 and dislocation absorption efficiency of $Z_{I}=1$ for SIA. Bulk concentrations are seen to be identical between the PF and DBC (1 nm mesh) implementation, and decrease with decrease in grain size. GB concentrations are higher in the PF implementation.
(a) 300°C
(b) 500°C
(c) 700°C
Figure S7: Evolution of bulk and GB concentrations of vacancy (vac.) and SIA in a 1D system of 1 $\mu$m grain size when under irradiation at $2\times 10^{-6}$ dpa/s at different temperatures. For each temperatures, plots for the RIS, RETS and RIS+RETS mechanisms are found to be nearly identical. Results for 500°C corresponds to Fig. 5.
(a) 300°C
(b) 500°C
(c) 700°C
Figure S8: Evolution of bulk and GB concentrations of Ni and Cr corresponding to Fig. S7. Both RIS and RETS are seen to evolve after the point defect concentrations in Fig. S7 reach a quasi-steady state determined by point defect generation, recombination, bulk dislocation absorption and GB sink. For 300°C and 700°C, RETS is seen to dominate, whereas for 500°C, RIS is seen to dominate. Atomic concentrations are seen to reach equilibrium RETS faster than steady-state RIS. Results for 500°C are identical to Fig. 5.
Figure S9: Steady-state point vacancy (vac.) and SIA concentrations in a 1D AM microstructure of 500 nm dislocation cell size irradiated at 500°C. The AM system corresponds to conditions similar to Fig. 10(a). Higher CW dislocation density $\rho_{b,w}$ results in significantly lower point defect concentrations at the CW. Biased absorption ($Z_{I}=1.2$) of SIA by dislocations results in a slightly lower SIA concentration.
Figure S10: Evolution of point defect concentrations at the CW and cell centers in the 1D AM microstructure of Fig. S9. Results are shown for different $Z_{I}$ and a CW dislocation density of $\rho_{b,w}=5\times 10^{15}$ m-2.
Figure S11: Point defect concentrations at (a) the dislocation cell center and (b) the dislocation CW center as a function of dislocation cell size from RIS simulations of 1D AM microstructure corresponding to Fig. 11. Increasing the dislocation density $\rho_{b,w}$ or the absorption efficiency for SIA $Z_{I}$ for $\rho_{b,w}=5\times 10^{15}$ m-2 is seen to decrease the point defect concentration at the CW. Significant decrease in the concentrations at the cell interior is only observed for low dislocation cell sizes.